\begin{thebibliography}{99}
	%\addcontentsline{toc}{chapter}{} 
	
	%	\bibitem{abra} Abramowitz, M.; Stegun, I. A. (1964) \textit{Handbook of mathematical
	%		functions with formulas, graphs, and mathematical tables. National Bureau of
	%		Standards Applied Mathematics Series, 55 For sale by the Superintendent of
	%		Documents}, U.S. Government Printing Office, Washington, D.C. 1964 xiv+1046
	%	pp. (Reviewer: D. H. Lehmer) 33.00 (65.05).
	%	
	
	%\bibitem{Adler e Taylor} Adler, R. J.; Taylor, J. E. (2007) Random fields and geometry. Springer Monographs in Mathematics. Springer, New York.
	
	\bibitem{AH12} Atkinson, K; Han, W. (2012) \textit{Spherical Harmonics and
		Approximations on the Unit Sphere: An introduction, Lecture Notes in
		Mathematics,} Volume 2044, Springer. 
	%	\bibitem{Azais e ws} Azais, J. M.; Wschebor, M. (2009) \textit{Level sets
	%		and extrema of random processes and fields}, Wiley and Sons, New Jersey.
	%	
	
	%\bibitem{BKMP2009} Baldi, P.; Kerkyacharian, G.; Marinucci, D.; Picard, D.
	%(2009). Asymptotics for Spherical Needlets, \textit{Annals of Statistics},
	%Vol. 37, No. 3, 1150-1171.
	
	%\bibitem{BKMP2009b} Baldi, P.; Kerkyacharian, G.; Marinucci, D.; Picard, D.
	%(2009). Subsampling Needlet Coefficients on the Sphere, \textit{Bernoulli},
	%Vol. 15, 438-463.
	
	\bibitem{BW} Beliaev, D.; Wigman, I. (2018).  Volume distribution of nodal
	domains of random band-limited functions. (English summary)  \textit{Probability Theory and Related Fields 172, no. 1-2, 453-492}. %	
	%	\bibitem{JW} Benatar, J.; Marinucci, D.; Wigman I. (2017) Planck-scale distribution of nodal length of arithmetic random waves, preprint,
	%	arXiv:1710.06153.
	%	
	%	\bibitem{Berry1977} Berry, M. V. (1977) Regular and irregular semiclassical wavefunctions. \textit{J. Phys.} A 10, no. 12, 2083?2091. 81.58.
	%	
	%	\bibitem{BW2016} Buckley, J.; Wigman, I. (2016)
	%	On the number of nodal domains of toral eigenfunctions.
	%	\textit{Ann. Henri PoincarÃ© 17}, no. 11, 3027-3062. 
	%	
	%	\bibitem{C} Cammarota, V. (2017) Nodal Area Distribution For Arithmetic Random
	%	Waves, arXiv:1708.07679v1.
	%	
	
	\bibitem{CM18} Cammarota, V.; Marinucci, D. (2018) A Quantitative Central
	Limit Theorem for the Euler-Poincar\'e Characteristic of Random Spherical
	Eigenfunctions. \textit{Annals of Probabability 46, no. 6, 3188-3228}.
	
	%\bibitem{CM15} Cammarota, V.; Marinucci, D. (2015) On the limiting behavior
	%of Needlets Polyspectra, (English, French version summary) \textit{\ Ann.
	%Inst. Henri Poincar\'e Prob. Stat.} 51, no. 3, 1159-1189.
	
	\bibitem{CMW2016F} Cammarota, V.; Marinucci, D.; Wigman, I. (2016)
	Fluctuations of the Euler-Poincar\'e characteristic for random spherical
	harmonics, \textit{Proceedings of the American Mathematical Society, no 11,
		4759-4775}.
	
	%\bibitem{CMW2016} Cammarota, V.; Marinucci, D.; Wigman, I. (2016) On the
	%distribution of the critical values of random spherical harmonics,  \textit{%
	%Journal of Geometric Analysis}, 4, 3252-3324. %	
	%	\bibitem{DNPR} Dalmao, F.; Nourdin, I.; Peccati, G.; Rossi, M. (2016) Phase Singularities in Complex	Arithmetic Random Waves. arXiv:1608.05631v3.
	%
	
	\bibitem{FXA} Feng, R.; Xu, X; Adler, R. J. (2018) Critical radius and
	Supremum of Random Spherical Harmonics (II). \textit{Electronic Communications in Probability 23, no. 50, 11 pp. 33C55}.
	%	\bibitem{11} Krishnapur, M.; Kurlberg P.; Wigman, I. (2013) Nodal length
	%	fluctuations for arithmetic random waves. \textit{Annals of Mathematics (2)
	%		177, no. 2, 699-737}.
	%	
	%\bibitem{LMX} Lan, X.; Marinucci, D.; Xiao, Y. (2018) Strong local nondeterminism and exact modulus of continuity for spherical Gaussian fields.
	%	\textit{Stochastic Processes and their Applications 128, no. 4, 1294-1315}. 
	
	%\bibitem{Lang e Schwab} Lang, A.; Schwab, C. (2015) Isotropic Gaussian random
	%	fields on the sphere: regularity, fast simulation and stochastic partial
	%	differential equations. \textit{Annals of Applied Probability 25, no. 6, 3047-3094}.
	
	%\bibitem{13} Lewis, A. (2012) The full squeezed CMB bispectrum from inflation \textit{Journal of Cosmology and Astroparticle Physics, 06, 023}. 
	%	
	%	\bibitem{M2008} Marinucci, D. (2008) A Central Limit Theorem and Higher Order Results for the Angular Bispectrum, \textit{Probability Theory and Related Fields}, Vol. 141, N.3-4, pp. 389-409, math.pr/0509430.
	%	
	%	\bibitem{M2006} Marinucci, D. (2006) High Resolution Asymptotics for the Angular Bispectrum of Spherical Random Fields, \textit{Annals of Statistics}, Vol.34, N.1, pp. 1-4.
	%	
	
	\bibitem{M e Peccati} Marinucci, D.; Peccati, G. (2011) \textit{Random
		fields on the sphere. Representation, limit theorems and cosmological
		applications}.  London Mathematical Society Lecture Note Series, 389.
	Cambridge University Press, Cambridge. 
	
	%\bibitem{MPRW} Marinucci, D.; Peccati, G.; Rossi, M.; Wigman, I. (2016)
	%Non-universality of nodal length distribution for arithmetic random waves 
	%\textit{Geom. Funct. Anal. 26}, no. 3, 926-960. 
	
	\bibitem{M e Mau 2015} Marinucci, D.; Rossi, M. (2015) Stein-Malliavin
	approximations for nonlinear functionals of random eigenfunctions on $%
	\mathbb{S}^d$. \textit{Journal of Functional Analysis 268, no. 8, 2379-2420}.
	
	\bibitem{MRW2017} Marinucci D.; Rossi M.; Wigman I. (2020+) The asymptotic
	equivalence of the sample trispectrum and the nodal length for random
	spherical harmonics. \textit{Annales de l'Institut Henri Poincar\'e,
		Probabilit\'es et Statistiques}, in press. 
	%	
	%	\bibitem{M e Vadlamani} Marinucci, D.; Vadlamani, S. (2015) A note on global suprema of band-limited spherical random functions. \textit{Statistics and Probability Letters 96, 141-148}.
	%	
	
	\bibitem{M e W 2012} Marinucci, D.; Wigman, I. (2014) On nonlinear
	functionals of random spherical eigenfunctions. \textit{Communications in Mathematical Physics
		327, no. 3, 849-872}. 
	
	\bibitem{M e W 2011} Marinucci, D.; Wigman, I. (2011) On the excursion sets
	of spherical Gaussian eigenfunctions. \textit{Journal of Mathematical Physics 52, no. 9,
		093301, 21 pp.}
	
	\bibitem{M e W 2011bis} Marinucci, D.; Wigman, I. (2011) The Defect Variance
	of Random Spherical Harmonics. \textit{Journal of Physics A: Mathematical
		and Theoretical, no 44, 355206}.
	
	%\bibitem{MVWBetal} McEwan, J. D.; Vielva, P.; Wiaux, Y.; Barreiro, R. B.; Cay%
	%ón, I.; Hobson, M P.; Lasenby, A. N.; Martínez-González, E. and Sanz, J. L.
	%(2007) Cosmological Applications of a Wavelet Analysis on the sphere. 
	%\textit{J. Fourier Anal. Appl. }13, no. 4,495-510. 
	%
	%\bibitem{21} Matsubara, T. (2010) Analytic Minkowski functionals of the
	%Cosmic  Microwave Background: second-order non-Gaussianity with bispectrum
	%and  trispectrum, \textit{Physical Review D, 81:083505}.
	
	%\bibitem{NPW2} Narcowich, F.J.; Petrushev, P.; Ward, J.D. (2006)
	%Decomposition of Besov and Triebel-Lizorkin spaces on the sphere, textit{%
	%Journal of Functional Analysis}, 238, 2, 530-564. 
	
	%\bibitem{NPW} Narcowich, F. J.; Petrushev, P.; Ward, J.D. (2006) Localized
	%tight frames on spheres, \textit{SIAM Journal of Mathematical Analysis}, 38,
	%2, 574-594.
	
	\bibitem{NS17} Nazarov, F.; Sodin, M. (2016). Asymptotic laws for the
	spatial distribution and the number of connected components of zero sets of
	Gaussian random functions. (English summary) \textit{Zhurnal Matematicheskoi Fiziki, Analiza, Geometrii 12, no. 3, 205-278}.
	
	\bibitem{NS09} Nazarov, F.; Sodin, M.(2009). On the number of nodal domains
	of random spherical harmonics. \textit{American Journal of Mathematics, no 131(5):1337-1357}. 
	
	\bibitem{Peccati Nourdin} Nourdin, I.; Peccati, G. (2012) \textit{Normal
		approximations with Malliavin calculus. From Stein's method to universality}.
	Cambridge Tracts in Mathematics, 192. Cambridge University Press, Cambridge. 
	%	
	%	\bibitem{NP2009} Nourdin, I.; Peccati, G. (2010) Stein's method meets Malliavin calculus: a short survey with new estimates. Recent development in stochastic dynamics and stochastic analysis, 207-236, \textit{Interdiscip. Math. Sci.}, 8, World Sci. Publ., Hackensack, NJ, 2010.
	%	
	%	\bibitem{P-R} Peccati, G.; Rossi, M. (2017) Nodal Statistics of Planar Random Waves, arXiv:1708.02281.
	%	
	%	\bibitem{N-P-R} Peccati, G.; Rossi, M. (2017). Quantitative limit theorems for local functionals of arithmetic random waves. \textit{Computation and Combinatorics in Dynamics, Stochastics and Control, The Abel Symposium 2016}, Springer (in press).  arXiv:1702.03765v1.
	%	
	
	%\bibitem{ORHMBBP} Oppizzi, F.; Renzi, A.; Liguori, M.; Hansen, F. K.;
	%Marinucci, D.; Baccigalupi, C.; Bertacca, D.; Poletti, D. (2019) Needlet
	%thresholding methods in component separation.
	%https://arxiv.org/abs/1911.01298
	
	%\bibitem{Rossi2} Rossi, M. (2015) On the high-energy behavior of nonlinear
	%functionals of random eigenfunctions on $\mathbb{S}^d$. \textit{Proceedings
	%of the 19th EYSM in Prague, 119-124}.
	%https://arxiv.org/abs/1506.01841.
	
	\bibitem{Rossi} Rossi, M. (2019) The defect of random hyperspherical
	harmonics. \textit{Journal of Theoretical Probability, no 32, 4, 2135-2165}.
	
	\bibitem{SW19} Sarnak, Peter; Wigman, Igor (2019) Topologies of nodal sets
	of random band-limited functions. \textit{Communications on Pure and Applied Mathematics 72, no. 2,
		275-342}.
	
	\bibitem{SW14} Sarnak, P.; Wigman, I. (2016) Topologies of nodal sets of
	random band limited functions. \textit{Advances in the theory of automorphic
		forms and their L-functions, 351-365, Contemp. Math., 664, Amer. Math. Soc.,
		Providence, RI.}
	
	%\bibitem{SMF10} Starck, J.-L.; Murtagh, F.; Fadili, J. (2010) \textit{Sparse
	%Image and Signal Processing: Wavelets, Curvelets, Morphological Diversity},
	%Cambridge University Press. 
	
	%	\bibitem{mau} Rossi, M. (2016) The geometry of spherical random fields, PhD thesis, arXiv:1603.07575.
	%	
	%	\bibitem{RW} Rudnick, Z.; Wigman I. (2016) Nodal intersections for random
	%	eigenfunctions on the torus, \textit{American Journal of Mathematics, 138, no. 6},
	%	1605-1644.
	%	
	%	%\bibitem{RW2014}
	%	%Rudnick, Z.; Wigman, I. Nodal intersections for random eigenfunctions on the torus. (English summary) 
	%	%\textit{Amer. J. Math. 138 } (2016), no. 6, 1605-1644. 
	%	
	%	\bibitem{RWY} Rudnick, Z.; Wigman, I.; Yesha N. (2015) Nodal intersections for
	%	random waves on the 3-dimensional torus, \textit{Annales Institut Fourier, 66}, no.
	%	6, 2455-2484.
	%	
	
	\bibitem{szego} Szego, G. (1975) Orthogonal polynomials, Fourth edition. 
	\textit{\ American Mathematical Society, Colloquium Publications, Vol. XXIII.%
	} \textit{American Mathematical Society, Providence, R.I.}
	
	\bibitem{Todino1} Todino, A. P. (2019) A Quantitative Central Limit Theorem
	for the Excursion Area of Random Spherical Harmonics over subdomains of $%
	\mathbb{S}^2$ \textit{Journal Mathematical Physics 60 , no. 2, 023505, 33
		pp.}
	
	\bibitem{Todino2} Todino, A. P. (2018) Nodal Lengths in Shrinking Domains
	for Random Eigenfunctions on $\mathbb{S}^2$. preprint. arXiv:1807.11787
	
	%	%\bibitem{Triebel} Triebel, H. (1995). \textit{Interpolation Theory, Function Spaces, Differential Operators}, 2nd ed. Johann Ambrosius Barth, Heildelberg. MR1328645.
	%	
	%	\bibitem{quantum theory} Varshalovich, D. A.; Moskalev, A. N.; Khersonski,
	%	V. K. (1988) \textit{Quantum theory of angular momentum. Irreducible tensors,
	%		spherical harmonics, vector coupling coefficients, 3nj symbols}. Translated
	%	from the Russian. World Scientific Publishing Co., Inc., Teaneck, NJ.
	%	
	
	\bibitem{W} Wigman, I. (2010) Fluctuations of the nodal length of random
	spherical  harmonics, \textit{Communications in Mathematical Physics, 398
		no. 3 787-831}.
	
	\bibitem{35} Wigman, I. (2009) On the distribution of the nodal sets of
	random  spherical harmonics. \textit{Journal of Mathematical Physics, 50,
		no. 1, 013521, 44 pp.}
\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[english]{babel}
\usepackage{fancyhdr}
\usepackage[latin1]{inputenc}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{latexsym}
\usepackage{amsthm}
\usepackage{color}
\usepackage{midpage}
\usepackage[width=15.00cm]{geometry}
\usepackage{amscd}
\usepackage{tikz-cd}
\usepackage{hyperref}

\numberwithin{equation}{section}
\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]
\theoremstyle{remark}
\newtheorem{rem}{Remark}[section]
\theoremstyle{cor}
\newtheorem{cor}[thm]{Corollary}
\theoremstyle{lem}
\newtheorem{lem}[thm]{Lemma}
\theoremstyle{prop}

\DeclareMathOperator{\Var}{Var}
\providecommand{\keywords}[1]{\textbf{Keywords and Phrases:} #1}
\providecommand{\classification}[1]{\textbf{AMS Classification:}#1}

\begin{document}

\author{{Anna Paola Todino} \\
{\small Ruhr-Universit\"at Bochum, Germany}\\
{\small anna.todino@ruhr-uni-bochum.de}}
\date{}
\title{Limiting Behavior for the Excursion Area of Band-Limited Spherical
Random Fields}
\maketitle

\begin{abstract}
In this paper, we investigate some geometric functionals for band-limited
Gaussian and isotropic spherical random fields in dimension 2. In
particular, we focus on the area of excursion sets, providing its behavior
in the high energy limit. Our result is based on Wiener chaos expansion for
non linear transform of Gaussian fields and on an explicit derivation on the
high-frequency limit of the covariance function of the field. As a simple corollary we establish also the Central Limit Theorem for the excursion area.

\end{abstract}


\begin{itemize} 
\item \keywords{} 
Gaussian Eigenfunctions, Excursion Area, Wiener-chaos expansion, Hilb's asymptotics, Central Limit Theorem.

\item \classification{} 60G60, 42C10, 33C55, 60F05.
\end{itemize}

\section{Introduction and Background}
Let $\{T_{\ell}(x),x\in \mathbb{S}^{2}\}$ denote the spherical harmonics, which are solutions of the Helmholtz equation: 
\begin{equation*}
\Delta _{\mathbb{S}^{2}}T_{\ell  }(x)+\ell (\ell +1)T_{\ell  }(x)=0,\mbox{ }\ell
=1,2,\dots;
\end{equation*}
where $\Delta _{\mathbb{S}^{2}}$ is the spherical Laplacian. We can put on these eigenfunctions a random structure such that $\{T_{\ell }(x),x\in \mathbb{S}^{2}\}$ are isotropic, centred Gaussian, with covariance function given by
\begin{equation*}
\mathbb{E}[T_{\ell }(x)T_{\ell }(y)]=\frac{2\ell +1}{4\pi }P_{\ell }(\cos
d(x,y)),
\end{equation*}
where $P_{\ell }$ is the Legendre polynomial and $d(x,y)$ the spherical
geodesic distance between $x$ and $y$, $d(x,y)=\arccos (\langle x,y\rangle )$.
After choosing a standard basis $\{Y_{\ell m}(x)\}$ of $L^2(\mathbb{S}^2)$, the random fields $\{T_{\ell }(x)\}$ can be expressed by
\begin{equation*}
T_{\ell }(x)=\sum_{m=-\ell }^{\ell }a_{\ell m}Y_{\ell m}(x),
\end{equation*}
where  $\{a_{\ell m}\}$ is the array of random spherical harmonic coefficients, which are independent, safe for the condition $\bar{a}_{\ell m} = (-1)^m a_{\ell,-m};$ for $m \ne0$ they are standard complex-valued Gaussian variables, while $a_{\ell 0}$ is a standard real-valued Gaussian variable. They satisfy
\begin{equation*}
\mathbb{E}[a_{\ell m}\bar{a}_{\ell ^{\prime }m^{\prime }}]=\delta _{\ell
}^{\ell ^{\prime }}\delta _{m}^{m^{\prime }}.
\end{equation*}
The geometry of the excursion sets of random eigenfunctions has been studied in many different papers, among them \cite{M e W 2012}, \cite{M e W 2011bis}, \cite{M e W 2011}, \cite{MRW2017}, \cite{W}, \cite{CMW2016F}, \cite{CM18}, \cite{35} and \cite{Todino1}, \cite{Todino2} for subdomains of $\mathbb{S}^2$ (see also \cite{M e Mau 2015} and \cite{Rossi} for the d-dimensional sphere $\mathbb{S}^d$). In this framework we aim to extend these results to the case of 
band-limited functions (see also \cite{BW},
\cite{NS09}, \cite{NS17}, \cite{SW19}, \cite{SW14}).\\

Hence let us consider the sequence $\alpha_{n,\beta}$ given by
\begin{equation}\label{alpha}
\alpha_{n,\beta}:= \sqrt{1-\frac{1}{n^\beta}}
\end{equation}
with $0 < \beta < 1, \beta \in \mathbb{R}.$ The band-limited functions here are random fields $\{\bar{T}_{\alpha_{n,\beta} }(x),x\in \mathbb{S}^{2}\}$ defined as 
\begin{equation}
\bar{T}_{\alpha_{n,\beta} }(x)=\sqrt{C_{n,\beta }}\sum_{\ell =\alpha_{n,\beta} n}^{n}T_{\ell
}(x),  \label{1}
\end{equation}%
where 
\begin{equation*}
C_{n,\beta }:=\frac{4\pi }{n^{2}(1-\alpha_{n,\beta} ^{2})+2n+1}= \frac{4\pi}{n^{2-\beta}+2n+1}.
\end{equation*}%
$\{\bar{T}_{\alpha_{n,\beta} }(x)\}$ are centred Gaussian with $\mathbb{E}[\bar{T}_{\alpha_{n,\beta} }(x)^{2}]=1$ and covariance
function given by 
\begin{equation}
\bar{\Gamma}_{\alpha_{n,\beta} }(x,y)= C_{n,\beta } \left(\sum_{\ell=\alpha_{n,\beta} n}^n \mathbb{E}[T_\ell(x)T_\ell(y)]\right)=C_{n,\beta }\sum_{\ell =\alpha_{n,\beta} n}^{n}\frac{%
	2\ell +1}{4\pi }P_{\ell }(\cos d(x,y)).  \label{cov}
\end{equation}%
We consider the excursion sets, defined as: 
\begin{equation}
A_{u}(\bar{T}_{\alpha_{n,\beta} }):=\{x\in \mathbb{S}^{2}:\bar{T}_{\alpha_{n,\beta} }(x)\geq u\},
\label{exsets}
\end{equation}%
with $u\in \mathbb{R}$, $u \ne 0$; in this paper we focus on the area of these regions, which we denote by $S_{\alpha_{n,\beta}}(u)$. Along the lines of \cite{M e W 2011}, it can be written as a functional of the random field itself in the following way
\begin{equation*}
S_{\alpha_{n,\beta}}(u)=\int_{\mathbb{S}^2} 1_{\{\bar{T}_{\alpha_{n,\beta}} (x) >u\}} \,dx,
\end{equation*}
where $1(\cdot)$ is, as usual, the characteristic function which takes value
one if the condition in the argument is satisfied, zero otherwise.  This
expression allows to project the area into the orthonormal system generated by
Hermite polynomials (Wiener chaoses projection) $H_{k}(u)$, $k \in \mathbb{N}$, that is
\begin{equation}
H_{0}(u)=1, H_{1}(u)=u, H_{2}(u)=u^{2}-1, \dots, H_{k}(t) = t H_{k-1}(t) - H^\prime_{k-1} (t ), k \geq 1.
\end{equation}
Indeed, since $1(\cdot) \in L^2(\mathbb{S}^2)$, it can
be expanded as
\begin{equation*}
1_{\{\bar{T}_{\alpha_{n,\beta}} (x) >u\}}=\sum_{q=0}^{\infty} \frac{J_q(u)}{q!} H_q(\bar{%
	T}_{\alpha_{n,\beta}}(x) ),
\end{equation*}
in $L^2(\Omega).$
The coefficients $\{J_q(\cdot)\}$ have the analytic expressions $%
J_0(u)=\Phi(u),$ $J_1(u)=-\phi(u),$ $J_2(u)=-u\phi(u),$ $J_3(u)=(1-u^2)%
\phi(u)$ and in general 
\begin{equation*}
J_q(u)=-H_{q-1}(u)\phi(u),
\end{equation*}
where $\phi(\cdot)$ and $\Phi(\cdot)$ are the density function and the
distribution function of a standard Gaussian variable (\cite{M e W 2011} and 
\cite{Peccati Nourdin}).
It follows that 
\begin{equation}\label{expansion}
S_{\alpha_{n,\beta}}(u)= \sum_{q=0}^{\infty} \dfrac{J_q(u)}{q!} \int_{\mathbb{S}^2} H_q(%
\bar{T}_{\alpha_{n,\beta}} (x) ) \, dx.
\end{equation}
Note that if we consider $\beta \equiv 1$, the random field in (\ref{1}) is the eigenfunction $T_\ell$ and it was shown in \cite{M e W 2011} that, in this case, 
the projection on the first component vanishes identically and the whole series in (\ref{expansion}) is dominated simply by the second chaotic component. More explicitly,
the variance of this single term is asymptotically equivalent to the variance of the full series, and its asymptotic distribution (Gaussian) gives also the limiting behavior of the excursion area. 
On the contrary, when $\beta \equiv 0$, the expansion in (\ref{expansion}) does not have any leading component, namely, each chaotic component has the same asymptotic behavior (as it happens
for the defect case, defined as the difference between positive and negative
regions, when only one eigenfunction is considered (see \cite{M e W
	2011bis}, \cite{Rossi})). 
It could
then be suspected that the limiting behavior may depend on the value of $%
\beta $, but this turns out not to be the case. Indeed, we will prove that for any $0<\beta< 1$ the second chaotic
component is still the leading term of the series expansion in (\ref{expansion}) and so no phase
transition with respect to $\beta$ arises.

\section{Main Result}
Let us consider the expansion of the excursion area given in (\ref{expansion}),
we can write
\begin{equation} \label{serie}
\begin{split}
S_{\alpha_{n,\beta}}(u) &= (1-\Phi(u) )\int_{\mathbb{S}^2}\,dx + \phi(u) \int_{\mathbb{S}%
	^2} H_1(\bar{T}_{\alpha_{n,\beta}} (x) ) \,dx\\&+ u\phi(u) \frac{1}{2} \int_{\mathbb{S}^2}
H_2(\bar{T}_{\alpha_{n,\beta}} (x) ) \,dx 
+ \sum_{q=3}^{\infty} \dfrac{J_q(u)}{q!} \int_{\mathbb{S}^2} H_q(\bar{T}
_{\alpha_{n,\beta}} (x) )\,dx,
\end{split}
\end{equation}
in the $L^2(\Omega)-$convergence sense.
Denoting 
\begin{equation*}
h_{{\beta};q}:=\int_{\mathbb{S}^2} H_q(\bar{T}_{\alpha_{n,\beta}} (x) ) \, dx \mbox{ }%
\mbox{ }\mbox{ }\mbox{ }\mbox{ }\mbox{ }\mbox{ } q=1,2,\dots,
\end{equation*}
we have that 
\begin{equation*}
S_{\alpha_{n,\beta}}(u)=\sum_{q=0}^{\infty} \dfrac{J_q(u)}{q!} h_{{\beta};q}.
\end{equation*}
\begin{rem}
	Note that $\int_{\mathbb{S}^2} H_1(\bar{T}_{\alpha_{n,\beta}} (x)) \,dx=0$, indeed 
	\begin{equation*}
	\int_{\mathbb{S}^2} H_1(\bar{T}_{\alpha_{n,\beta}} (x)) \,dx= \int_{\mathbb{S}^2} \bar{T}%
	_{\alpha_{n,\beta}} (x) \,dx= \sqrt{C_{n,{\beta}}} \sum_{\ell={\alpha_{n,\beta}} n}^{n} \int_{\mathbb{%
			S}^2} T_\ell(x) \, dx=0
	\end{equation*}
	thanks to the orthogonal property of Spherical Harmonics (see \cite{M e Peccati}, page 66).
\end{rem}
\begin{rem}
	The choice of $C_{n,{\beta}}$ in (\ref{cov}) is such that $%
	\Var[\bar{T}_{\alpha_{n,\beta}}(x)]=1$. Indeed,  
	\begin{equation*}  \label{key}
	\begin{split}
	\Var&[{\bar{T}_{\alpha_{n,\beta}}(x)}]=\!\dfrac{4\pi}{n^2(1-{\alpha^2_{n,\beta}})+2n+1}
	\!\!\sum_{\ell={\alpha_{n,\beta}} n}^{n}\!\!\![\Var{T_\ell(x)}]=\dfrac{4\pi}{n^2(1-{\alpha^2_{n,\beta}})+2n+1}\!\!\!
	\sum_{\ell={\alpha_{n,\beta}} n}^{n} \!\!\!\dfrac{2\ell+1}{4\pi} \\
	&=\dfrac{1}{n^2(1-{\alpha^2_{n,\beta}})+2n+1} \bigg[ 2\bigg(\dfrac{n(n+1)}{2} - \dfrac{%
		({\alpha_{n,\beta}} n-1)({\alpha_{n,\beta}} n)}{2}\bigg) +( n-{\alpha_{n,\beta}} n+1)\bigg] \\
	&= \dfrac{1}{n^2(1-{\alpha^2_{n,\beta}})+2n+1} ( n^2+n -{\alpha^2_{n,\beta}} n^2+{\alpha_{n,\beta}} n+n -{\alpha_{n,\beta}}
	n +1)=1.
	\end{split}%
	\end{equation*}
\end{rem}
The main result of this paper gives
the high energy behavior of the variances of $h_{\beta; q}$ for $q\geq2$.
\begin{thm}
	\label{main} 	For $0 < \beta < 1,$
	as $n \rightarrow \infty$, 
	\begin{align}
	\Var(h_{\beta; 2})&=%
	\dfrac{32\pi^2}{n^{2-\beta}}- \dfrac{64\pi^2}{n^{3-2\beta}} +o\left(\frac{1}{n^{3-2\beta}}\right) & \mbox{ for } q=2, \\ 
	\Var(h_{\beta;q})& = 
	O\left( \dfrac{1}{n^{2}}\right) & \mbox{  for  } q
	= 3 \mbox{ and } q \geq 5 \\
	\Var(h_{\beta;4})&=O\left( \dfrac{ \log n}{n^{2}}\right) & \mbox{  for  } q
	= 4.
	\end{align}
\end{thm}
\begin{rem}From Theorem \ref{main} it follows that the second chaos is the leading term for all $\beta \in (0,1)$. 
	If $\beta =1$, the random field in (\ref{1}) is simply one random eigenfunction and this case has already been investigated in \cite{M e W 2011}, where it has been proved that the variance of $h_{1;2}$ is $32\pi^2\dfrac{2}{2\ell+1}$. 
\end{rem}
For the continuity of the norm and the orthogonality of the
Hermite polynomials,
the following expansion holds in the $L^2(\Omega)$ sense: 
\begin{equation*}  \label{serie-var}
\mbox{Var}\lbrack\ \!S_{\alpha_{n,\beta}}(u) ] \!= \!0 +\! 0+\! \dfrac{u^2\phi(u)^2}{4} 
\mbox{Var}\bigg[ \int_{\mathbb{S}^2} H_2(\bar{T}_{\alpha_{n,\beta}} (x)) \,dx %
\bigg] + \sum_{q=3}^{\infty} \dfrac{J_q(u)^2}{q!^2} \mbox{Var}\bigg[ \!\int_{%
	\mathbb{S}^2} \!H_q(\bar{T}_{\alpha_{n,\beta}} (x)) \,dx\bigg].
\end{equation*}
Then, as a corollary, we get
\begin{cor} 
	\label{corollary} For $0 < \beta <1 $, as $n \rightarrow \infty,$  
	\begin{equation*}
	\Var(S_{\alpha_{n,\beta}}(u))= \dfrac{u^2\phi(u)^2}{4} \Var(h_{{\beta};2})+O \left(\frac{\log n}{n^{2}}\right)=32 \pi^2  \dfrac{u^2\phi(u)^2}{4} \dfrac{1}{n^{2-\beta}}+ R_\beta
	\end{equation*}
	where  
	\begin{equation}
	R_\beta= 
	\begin{cases}
	O\left(\dfrac{\log n}{n^2}\right) & \mbox{ if }  0 < \beta \leq 1/2, \\ 
	O\left(\dfrac{1}{n^{3-2\beta}}\right)  & \mbox{ if  }  \frac{1}{2}<\beta<1. \\ 
	\end{cases}%
	\end{equation}
\end{cor}
The key role of the proof of Theorem \ref{main} is played by the derivation
of the asymptotic behaviour for the covariance function in (\ref{cov}) and
it is given in Lemma \ref{Lemma1}.\\

Denoting by $N$ the North Pole, we fix $x=N$ and, in view of the isotropy,
we can write $\bar{\Gamma}_{\alpha_{n,\beta}}(x,y)=\bar{\Gamma}_{\alpha_{n,\beta}}(\cos
\theta)$ with $\theta \in [0,\pi)$. Changing variable $\theta= \frac{\psi}{\alpha_{n,\beta} n}$,  the following lemma gives the
asymptotic behaviour in the high frequency limit of the covariance function
for $1<\psi \leq \alpha_{n,\beta} n (\pi-\epsilon)$, for any $\epsilon>0$.
\begin{lem}
	\label{Lemma1}  Given $\bar{\Gamma}_{\alpha_{n,\beta}}(x,y)$ as in (\ref{cov}%
	), for $0 < \beta <1,
	\beta \in \mathbb{R},$ we have that, for $1<\psi < n^{\beta}$, as $n \rightarrow \infty$,  
	\begin{equation*}  \label{asymptotic cov function}
	\begin{split}
	\bar{\Gamma}_{\alpha_{n,\beta}}\left(\cos \frac{\psi}{\alpha_{n,\beta} n}\right)
	&=\frac{C_{n,\beta}}{4\pi}  \frac{(\alpha_{n,\beta} n)^2}{\sqrt{\psi}}   \frac{1}{n^\beta} \sqrt{\frac{2}{\pi}} \frac{1}{2} \bigg(\!\!-\cos \left(\psi-\frac{3\pi}{4}\right) \frac{\psi}{4n^{\beta}}- \sin  \left(\psi-\frac{3\pi}{4}\right) \\&+O\left( \frac{1}{ \psi} \right)\bigg)=:\bar{\Gamma}_1;
	\end{split}
	\end{equation*}
	while, for $\psi \in [n^{\beta},\alpha_{n,\beta} n (\pi-\epsilon)]$ it is 
	\begin{equation*}\label{cov2}
	\begin{split}
	\bar{\Gamma}_{\alpha_{n,\beta}}&\left(\cos \frac{\psi}{\alpha_{n,\beta} n}\right) =\frac{C_{n,\beta}}{4\pi} 2
	\sqrt{\frac{2}{\pi}}\left( -2 \sin \left(\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi\right) \sin\left(\frac{n+1}{2n\alpha_{n,\beta}}\psi-\frac{\psi}{2}\right)\right) \\&\times\left(\frac{(\alpha_{n,\beta} n )^2}{\psi^{3/2}} +  \sqrt{\psi} \right)
	+O\left(C_{n,\beta}(\alpha_{n,\beta}n)^2 \frac{1}{\psi^{5/2}}\right) =:\bar{\Gamma}_2.
	\end{split}
	\end{equation*}
\end{lem}
Moreover, we observe that 
\begin{equation}\label{CLT}
\begin{split}
\int_{\mathbb{S}^2} H_2(\bar{T}_{\alpha_{n,\beta}}(x)) \, dx &= \int_{\mathbb{S}^2} (\bar{T}_{\alpha_{n,\beta}}(x)^2-1) \, dx= C_{n,\beta} \sum_{\ell}\sum_{\ell'} \int_{\mathbb{S}^2} T_\ell(x) T_{\ell'}(x) \, dx -4\pi \\&= C_{n,\beta} \sum_{\ell=\alpha_{n,\beta} n}^{n}\sum_{m=-\ell}^{\ell} |a_{\ell m}|^2 -4\pi ,
\end{split}
\end{equation}
which are sums of independent Gaussian random variables; the mean of (\ref{CLT}) is zero and then, from Theorem \ref{main}, the Central Limit Theorem follows for $h_{\beta;2}$.
As a consequence, along the same lines as in \cite{M e Mau 2015}, we can establish the validity of the Central Limit Theorem for $S_{\alpha_{n,\beta}}$. Note that from (\ref{serie}) it is easy to see that $\mathbb{E}[S_{\alpha_{n,\beta}} (u)]= (1-\Phi(u))4\pi$.
\begin{cor}\label{TLC}
	For all $0 <\beta < 1$, as $n \rightarrow \infty$, $$\dfrac{S_{\alpha_{n,\beta}} (u)-\mathbb{E}[S_{\alpha_{n,\beta}} (u)]}{\sqrt{\Var[S_{\alpha_{n,\beta}} (u)]}} \rightarrow_d Z,$$
	where $Z \sim N(0,1)$ and $\rightarrow_d $ denote convergence in distribution.
\end{cor}

\begin{rem}
	When $\beta=0$ all the chaotic components have the same asymptotic behavior. Indeed, the idea is that the covariance function in this case behaves like
	\begin{equation}
	\bar{\Gamma}_{\alpha_{n,0}}\left(\cos \frac{\psi}{n}\right) \sim C_{n,0} \frac{1}{\sin (\psi/n)} \left(\frac{\psi/n}{\sin (\psi/n)}\right)^{1/2} (n+1) \sqrt{\frac{2}{(n+1)(\psi/n)}} \cos \left(\psi- \frac{3\pi}{4}\right)
	\end{equation} so that for all $q$  $$\Var(h_{0,q}) \sim \frac{C_q}{n^2} \int_{1}^{\pi n} \frac{1}{\psi^{\frac{3}{2}q-1}} \cos \left(\psi-\frac{3\pi}{4}\right)^q \,d\psi\sim \frac{C'_q}{n^2}$$ since the integral converges.
\end{rem}

\section{Proof of the Main Result (Theorem \protect\ref{main})}
\begin{proof}[Proof of Theorem \ref{main} assuming Lemma \ref{Lemma1}]
	First of all we remind the following
	property (see for instance \cite{M e Peccati}, page 98): let $Z_1, Z_2$ be jointly Gaussian; then, for all $q_1,
	q_2 \geq 0$ 
	\begin{equation}  \label{propertyH}
	\mathbb{E}[H_{q_1}(Z_1)H_{q_2}(Z_2)] = q_1!\delta_{q_1}^{q_2} \mathbb{E}[Z_1Z_2].
	\end{equation}
	Now we start computing the variance of $h_{\beta;2}$; hence,
	\begin{equation}\label{2}
	\begin{split}
	& \Var(h_{\beta;2})=\Var \bigg[ \int_{\mathbb{S}^2} H_2(\bar{T}_{{\alpha_{n,\beta}}}(x)) \, dx\bigg]=\mathbb{E} \bigg[ \int_{\mathbb{S}^2} H_2(\bar{T}_{{\alpha_{n,\beta}}}(x)) \, dx  \bigg]^2\\&= \mathbb{E} \bigg[ \int_{\mathbb{S}^2 \times\mathbb{S}^2  }\!\!\!\! H_2(\bar{T}_{{\alpha_{n,\beta}}}(x)) H_2(\bar{T}_{{\alpha_{n,\beta}}}(y))\, dx dy \bigg]=\int_{\mathbb{S}^2\times \mathbb{S}^2} \!\!\!\! \mathbb{E}[H_2(\bar{T}_{{\alpha_{n,\beta}}}(x)) H_2(\bar{T}_{{\alpha_{n,\beta}}}(y))] \,dx dy
	\end{split}
	\end{equation}
	which is, in view of (\ref{propertyH}), equal to
	$$\int_{\mathbb{S}^2  \times \mathbb{S}^2} 2\bar{\Gamma}_{\alpha_{n,\beta}}(x,y)^2 \,dx dy.$$ Using (\ref{cov}) we get
	\begin{equation}\label{3.3}
	\begin{split}
	\Var(h_{\beta;2})=2 C_{n,\beta}^2 \int_{\mathbb{S}^2  \times \mathbb{S}^2}\sum_{\ell={\alpha_{n,\beta}}n}^{n} \sum_{\ell'={\alpha_{n,\beta}} n}^{n}\frac{2\ell+1}{4\pi} P_\ell(\langle x,y \rangle) \frac{2\ell'+1}{4\pi} P_{\ell'}(\langle x,y \rangle) \, dx dy
	\end{split}
	\end{equation}
	and exchaging integrals and sums, and 
	applying the duplication property (see \cite{M e Peccati}, Ch. 3), that is,
	$$ \int_{ \mathbb{S}^2} \dfrac{2\ell+1}{4\pi} P_\ell(\langle x,y \rangle) \dfrac{2\ell^\prime+1}{4\pi} P_{\ell^\prime}(\langle y,z \rangle) \,dy = \dfrac{2\ell+1}{4\pi} P_\ell(\langle x,z \rangle) \delta_{\ell}^{\ell^\prime},$$
	(\ref{3.3}) is equal to
	$$ 2 C_{n,\beta}^2 \sum_{\ell} \frac{2\ell+1}{4\pi} \int_{ \mathbb{S}^2} P_\ell(\langle x,x \rangle)\,dx$$
	and since $P_\ell(0)=1 \mbox{ } \mbox{ }\forall \ell$, we conclude that
	\begin{equation*}
	\begin{split}
	\Var(h_{\beta;2})  &=2 C_{n,\beta}^2 \sum_{\ell=\alpha_{n,\beta} n}^{n} \frac{2\ell+1}{4\pi} \int_{ \mathbb{S}^2} \,dx=2 C_{n,\beta}^2 \sum_{\ell=\alpha_{n,\beta} n}^{n} \frac{2\ell+1}{4\pi} 4\pi \\& =2 C_{n,\beta}^2 [( n(n+1)-\alpha_{n,\beta} n(\alpha_{n,\beta} n-1))+n+1-\alpha_{n,\beta} n ]
	\\&= 2 C_{n,\beta}^2(n^2(1-\alpha_{n,\beta}^2)+2n+1) =\dfrac{2 (4\pi)^2}{n^2(1-\alpha_{n}^2)+2n+1}= \dfrac{2(4\pi)^2}{n^{2-\beta}+2n+1}.
	\end{split}
	\end{equation*}
	Finally, Taylor expansion implies
	\begin{equation}\label{C2}
	\begin{split}
	\Var(h_{\beta;2})=& \dfrac{32\pi^2}{n^{2-\beta}} \left(  1-\frac{2}{n^{1-\beta}}+o\left(\frac{1}{n^{1-\beta}}\right)\right)= \dfrac{32\pi^2}{n^{2-\beta}}- \dfrac{64\pi^2}{n^{3-2\beta}} +o\left(\frac{1}{n^{3-2\beta}}\right).
	\end{split}
	\end{equation}
	Let us focus now on the variance of the chaotic components $h_{\beta;q}$, for $q>2$. Hence, same computations as in (\ref{2}) lead to
	\begin{equation*}\label{eq}
	\begin{split}
	\Var(h_{\beta;q})&=\Var\bigg(\int_{ \mathbb{S}^2}H_q(\bar{T}_{\alpha_{n,\beta}}(x))\,dx\bigg)=\int_{\mathbb{S}^2 \times \mathbb{S}^2}\mathbb{E}[H_q(\bar{T}_{\alpha_{n,\beta}}(x))H_{q'}(\bar{T}_{\alpha_{n,\beta}}(y))]\, dx dy
	\\&	=q!\int_{\mathbb{S}^2  \times \mathbb{S}^2} \bar{\Gamma}_{\alpha_{n,\beta}}(x,y)^q \,dx dy
	\end{split}
	\end{equation*}
	and for the isotropy it is
	\begin{equation}\label{integ}
	\begin{split}
	&	= 2\pi |\mathbb{S}^2|q!\int_{0}^{\pi} \bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)^q \sin \theta d\theta= 2\pi |\mathbb{S}^2|q!\int_{0}^{1/n} \bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)^q \sin \theta d\theta\\&+2\pi |\mathbb{S}^2 |q!\int_{1/n}^{1/{(\alpha_{n,\beta} n)}} \bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)^q \sin \theta d\theta +
	2\pi |\mathbb{S}^2 |q!\int_{1/{(\alpha_{n,\beta} n)}}^{\pi} \bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)^q \sin \theta d\theta.
	\end{split}
	\end{equation}
	Note that for $ \theta \in [0,1/n]$, since $| \bar{\Gamma}_{\alpha_{n,\beta}}(x,y)| \leq 1$, changing variable $\theta=\dfrac{\psi}{n}$, we have that 
	\begin{equation}\label{2var}
	\int_{0}^{1/n} \bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)^q \sin \theta \,d\theta =O \left( \int_{0}^{1/n} |\sin \theta |\, d\theta \right)= O \left( \frac{1}{n} \int_{0}^{1} \frac{\psi}{n} \,d\psi \right)= O \left(\frac{1}{n^{2}} \right).
	\end{equation}
	In the same way
	\begin{equation}\label{3var}
	\begin{split}
	\int_{1/n}^{1/(\alpha_{n,\beta} n)} &\bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)^q \sin \theta \,d\theta =O \left( \int_{1/n}^{1/(\alpha_{n,\beta} n)} |\sin \theta |\, d\theta \right)= O \left( \frac{1}{n} \int_{1}^{1/\alpha_{n,\beta}} \frac{\psi}{n} \,d\psi \right)\\&= O \left(\frac{1}{2n^{2}} \left(\frac{1}{\alpha_{n,\beta}^2}-1\right) \right) =O \left(\frac{1}{2n^{2}} \left(1+\frac{1}{n^\beta} -1\right)\right)=O\left( \frac{1}{n^{2+\beta}} \right)
	\end{split}
	\end{equation}
	where we have used the geometric series for $\frac{1}{\alpha_{n,\beta}^2}=\frac{1}{1-1/n^\beta}$.
	For the last integral in (\ref{integ}), changing the variables as $\theta= \frac{\psi}{\alpha_n n}$, we have that
	\begin{equation}
	\begin{split}
	&\int_{1/(\alpha_{n,\beta} n)}^{\pi} \bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)^q \sin \theta d\theta = 
	\frac{1}{\alpha_{n,\beta} n}\int_{1}^{\pi\alpha_{n,\beta} n} \bar{\Gamma}_{\alpha_{n,\beta}}\left(\cos \frac{\psi}{\alpha_{n,\beta} n}\right)^q\sin \left(\frac{\psi}{\alpha_{n,\beta} n}\right)d\psi\\&
	=\frac{1}{\alpha_{n,\beta} n}\int_{1}^{n^{\beta}} \bar{\Gamma}_1^q\sin \left(\frac{\psi}{\alpha_{n,\beta} n}\right)d\psi+\frac{1}{{\alpha_{n,\beta} n}}\int_{n^{\beta}}^{\pi\alpha_{n,\beta} n} \bar{\Gamma}_{2}^q\sin \left(\frac{\psi}{\alpha_{n,\beta} n}\right)d\psi=:I_1+I_2.
	\end{split}
	\end{equation}
	From Lemma \ref{Lemma1}, taking the $q-$power, we see that
	\begin{equation}\label{gammaq1}
	\begin{split}
	\Gamma_1^q&=\frac{C_{n,\beta}^q}{(4\pi)^q}  \frac{(\alpha_{n,\beta} n)^{2q}}{\psi^{q/2}} \frac{1}{2^q}\sqrt{\frac{2}{\pi}}^q  \frac{1}{n^{\beta q}} \bigg(-\cos \left(\psi-\frac{3\pi}{4}\right) \frac{\psi}{4n^{\beta}}- \sin  \left(\psi-\frac{3\pi}{4}\right) \bigg)^q\\&+ O \left( C_{n,\beta}^q \frac{(\alpha_{n,\beta} n)^{2q}}{n^{\beta q}} \frac{1}{\psi^{q/2+1}} \bigg(-\cos \left(\psi-\frac{3\pi}{4}\right) \frac{\psi}{4n^{2\beta}}- \sin  \left(\psi-\frac{3\pi}{4}\right)\bigg)^{q-1}\right)
	\end{split}%
	\end{equation}
	and 
	\begin{equation}\label{gammaq2}
	\begin{split}
	\Gamma_2^q&=\frac{C_{n,\beta}^q}{(4\pi)^q} 2^q \left(\frac{2}{\pi}\right)^{q/2}(\alpha_{n,\beta} n )^{2q}
	\sqrt{\psi}^q (-2)^q
	\sin \left(\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)\psi}{2\alpha_{n,\beta} n}\right)^q \sin\left(\frac{(n+1)\psi}{2n\alpha_{n,\beta}}-\frac{\psi}{2}\right)^q \\& \left[ \frac{1}{\psi^2}+\frac{1}{(\alpha_{n,\beta}n)^2}\right]^q
	+O\left(C_{n,\beta}^q   \frac{(\alpha_{n,\beta} n )^{2q}}{\psi^{-q/2+2}} \sin \left(\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi\right)^{q-1} \right. \\&\left.\sin\left(\frac{n+1}{2n\alpha_{n,\beta}}\psi-\frac{\psi}{2}\right)^{q-1}\times\left[ \frac{1}{\psi^2}+\frac{1}{(\alpha_{n,\beta}n)^2}\right]^{q-1} \right).
	\end{split}
	\end{equation}
	As far as $I_2$ is concerned, using the Taylor expansion of $\sin \psi/(\alpha_{n,\beta} n)$, the Newton's binomial formula and (\ref{gammaq2}) we get
	\begin{equation}
	\begin{split}
	&I_2=\frac{1}{\alpha_{n,\beta} n } \int_{n^\beta}^{n\alpha_{n,\beta}\pi} \bar{\Gamma}_2^q \sin\frac{\psi}{\alpha_{n,\beta}} \,d\psi=\frac{1}{(\alpha_{n,\beta}n)^2}\int_{n^\beta}^{n\alpha_{n,\beta}\pi}\bar{\Gamma}_2^q \psi\,d\psi\\&+O\left(\frac{1}{(\alpha_{n,\beta}n)^4}\int_{n^\beta}^{n\alpha_{n,\beta}\pi}\bar{\Gamma}_2^q \psi^3\,d\psi\right)= \left(\frac{C_{n,\beta}}{4\pi} \right)^q (2)^q \left(\frac{2}{\pi}\right)^{\frac{q}{2}} (-2)^q \frac{(\alpha_{n,\beta} n)^{2q}}{(\alpha_{n,\beta}n)^2}
	\int_{n^\beta}^{\alpha_{n,\beta} n \pi-\epsilon} \!\!\! \sqrt{\psi}^q \\&\times \sum_{k=0}^{q} \binom{q}{k} \frac{1}{\psi^{2k}} \frac{1}{(\alpha_{n,\beta}n)^{2(q-k)}} \psi \,d\psi
	\times \left(\sin \left(\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi\right) \sin\left(\frac{n+1}{2n\alpha_{n,\beta}}\psi-\frac{\psi}{2}\right)\right)^q \,d\psi\\&
	+ O\left( \frac{1}{(\alpha_{n,\beta} n )^2}  \int_{n^{\beta}}^{\pi n \alpha_{n,\beta}-\epsilon} C_{n,\beta}^q   \frac{(\alpha_{n,\beta} n )^{2q}}{\psi^{-q/2+1}} \sin \left(\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi\right)^{q-1}\right.\\& \times \left. \sin\left(\frac{n+1}{2n\alpha_{n,\beta}}\psi-\frac{\psi}{2}\right)^{q-1}\left[ \frac{1}{\psi^2}+\frac{1}{(\alpha_{n,\beta}n)^2}\right]^{q-1} \,d\psi \right)+ O\left( \frac{1}{(\alpha_{n,\beta}n)^{4}} \int_{n^{\beta}}^{\pi n \alpha_{n,\beta}} \Gamma_2^q \psi^{3} \,d\psi\right).
	\end{split}
	\end{equation}
	Let us focus on $ (\alpha_{n,\beta} n)^{2q-2}\frac{1}{(n^{2-\beta}+2n+1)^q}$. Taylor expansion implies that
	\begin{equation}\label{Taylorexp}
	\begin{split}
	\frac{n^{2q-2}}{(n^{2-\beta})^q} \frac{1}{\left( 1+\frac{2n+1}{n^{2-\beta}}\right) ^q} \left(1-\frac{1}{n^{\beta}}\right)^{q-1}&= \frac{n^{-2}}{n^{-\beta q}} \left(1+O\left(\frac{2n+1}{n^{2-\beta}}\right)\right) \left( 1+O\left(\frac{1}{n^{\beta}}\right)\right) \\&=\frac{n^{-2}}{n^{-\beta q}} \left(1+O\left(\frac{1}{n^{\beta}}\right)+O\left(\frac{1}{n^{1-\beta}}\right)\right).
	\end{split}
	\end{equation}
	Hence for $q \geq 4$, bounding the sine by 1 and solving the integral in $\psi$, we can see that $$I_2\ll C_q\frac{n^{q\beta}}{n^2}  \frac{1}{(\alpha_{n,\beta} n)^{2q}} (\alpha_{n,\beta}n)^{q/2+2} =O\left(\frac{1}{n^2}\right);$$
	while for $q=3$, we have to compute
	\begin{equation}\label{q3}
	\begin{split}
	C_3 \frac{n^{3\beta}}{n^2} &\int_{n^\beta}^{\alpha_{n,\beta} n \pi} \psi^{3/2+1}
	\left( \frac{1}{\psi^6}+ \frac{1}{\psi^4 (\alpha_{n,\beta}n)^2}+ \frac{1}{\psi^2 (\alpha_{n,\beta}n)^4}+\frac{1}{(\alpha_{n,\beta}n)^6} \right)\\&
	\times \sin \left(\psi-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi \right)^3 \sin\left(\frac{n+1}{2\alpha_{n,\beta}n}\psi-\psi\right)^3 \,d\psi .
	\end{split}
	\end{equation}
	In view of the following formula
	\begin{equation*}
	\begin{split}
	\sin(a)^3\sin(b)^3&= \frac{1}{32} \bigg[-3 \cos(a-3b)+\cos(3a-3b)+9\cos(a-b)-3\cos(3a-b)\\&-9 \cos (a+b)+3\cos(3a+b)+3\cos(a+3b)-\cos(3a+3b)\bigg],
	\end{split}
	\end{equation*}
	integration by parts gives the convergence of the integral in (\ref{q3}) multiplied by $n^{3\beta}$ so that (\ref{q3}) is $O(1/n^2)$.
	Regarding $I_1$ applying again the Taylor expansion of $\sin\psi/(\alpha_{n,\beta}n)$ and using (\ref{gammaq1}) we get
	\begin{equation}\label{perq3}
	\begin{split}
	&I_1=\frac{1}{\alpha_{n,\beta} n } \int_{1}^{n^\beta} \Gamma_1^q \sin\frac{\psi}{\alpha_{n,\beta}} \,d\psi=\frac{1}{(\alpha_{n,\beta}n)^2}\int_{1}^{n^\beta} \Gamma_1^q \psi\,d\psi+O\left(\frac{1}{(\alpha_{n,\beta}n)^4}\int_{1}^{n^\beta} \Gamma_1^q \psi^3\,d\psi\right)\\&
	= \frac{C_{n,\beta}^q}{(4\pi)^q} \frac{1}{2^q}\left(\frac{2}{\pi}\right)^{q/2}\!\!\! \frac{1}{n^{\beta q}}(\alpha_{n,\beta} n)^{2q-2}  \int_{1}^{n^\beta} \frac{1}{\psi^{q/2}}\psi \bigg(-\cos \left(\psi-\frac{3\pi}{4}\right) \frac{\psi}{4n^{\beta}}- \!\!\sin  \left(\psi-\frac{3\pi}{4}\right) \bigg)^q\!\! \,d\psi\\&+ O \left( C_{n,\beta}^q \frac{(\alpha_{n,\beta} n)^{2q-2}}{n^{\beta q}} \int_{1}^{n^\beta}\frac{1}{\psi^{q/2}} \bigg(-\cos \left(\psi-\frac{3\pi}{4}\right) \frac{\psi}{4n^{2\beta}}- \sin  \left(\psi-\frac{3\pi}{4}\right)\bigg)^{q-1}\,d\psi\right).
	\end{split}
	\end{equation}
	Then, bounding $\bigg|\left(-\cos\left(\psi-\frac{3\pi}{4}\right) \frac{\psi}{4n^{\beta}}- \sin  \left(\psi-\frac{3\pi}{4}\right)\right)^q\bigg|\leq |(\frac{\psi}{4n^{\beta}}+1)^q|$ and using the Newton's binomial formula, we can deduce that, for all $q\geq 4$,
	\begin{equation}
	\begin{split}
	I_1&\sim\frac{C_{n,\beta}^q}{(4\pi)^q} \frac{1}{2^q}\left(\frac{2}{\pi}\right)^{q/2} \frac{1}{n^{\beta q}}(\alpha_{n,\beta} n)^{2q-2} \int_{1}^{n^\beta} \frac{1}{\psi^{q/2-1}}
	\sum_{k=0}^{q} \binom{q}{k} \left( \frac{\psi}{4n^\beta}\right)^k  \,d\psi\\&
	= \frac{C_{n,\beta}^q}{(4\pi)^q} \frac{1}{2^q}\left(\frac{2}{\pi}\right)^{q/2}\frac{1}{n^{\beta q}}(\alpha_{n,\beta} n)^{2q-2}\sum_{k=0}^{q} \binom{q}{k} \left(\frac{1}{4n^\beta}\right)^k \int_{1}^{n^\beta} \psi^{1-q/2+k}  \,d\psi
	\end{split}
	\end{equation}
	and (\ref{Taylorexp}) gives
	\begin{equation*}
	\begin{split}
	I_1&=O\bigg( \frac{1}{n^2} \bigg(\sum_{k=0, k\ne q/2-1,k\ne q/2-2}^{q} \binom{q}{k} \left(\frac{1}{4n^\beta}\right)^k  \psi^{2 -\frac{q}{2}+k}\bigg|_{1}^{n^\beta} +\binom{q}{\frac{q}{2}-1} \left(\frac{1}{4n^\beta}\right)^{\frac{q}{2}-1}(n^\beta-1)\\&+ \binom{q}{q/2-2} \left(\frac{1}{4n^\beta}\right)^{q/2-2} \int_{1}^{n^\beta} \psi^{-1} \,d\psi\bigg) \bigg)=O\left(\frac{1}{n^2}\frac{1}{n^{q\beta/2-2\beta}}\right)+O\left(\frac{1}{n^2}\frac{\log n}{n^{q\beta/2-2\beta}}\right)
	\end{split}
	\end{equation*}
	Note that for $q=4$ we obtain $I_1=O(\log n /n^2)$.
	It remains to study the case $q=3$. Developping the third power in (\ref{perq3}), and in view of (\ref{Taylorexp}) we get
	\begin{equation*}
	\begin{split}
	&
	I_1 \sim \frac{C_q}{n^2}
	\int_{1}^{n^{\beta}} \frac{1}{\sqrt{\psi}} \bigg(\frac{\psi^{3}}{4n^{3\beta}}  \cos \left(\psi -\frac{3\pi}{4}\right)^3- \sin\left(\psi- \frac{3\pi}{4}\right)^{3}\\&- \frac{3\psi^{2}}{16n^{2\beta}}  \cos \left(\psi -\frac{3\pi}{4}\right)^2 \sin\left(\psi- \frac{3\pi}{4}\right)- \frac{3\psi}{4n^{\beta}}  \cos \left(\psi -\frac{3\pi}{4}\right)\sin\left(\psi- \frac{3\pi}{4}\right)^{2}\bigg)\,d\psi.
	\end{split}
	\end{equation*}
	Integrating by parts it can be seen that the integral converges. For example,
	\begin{equation*}
	\begin{split}
	&\int_{1}^{n^\beta} \!\!\!\frac{\sin \left(\psi - \frac{3 \pi}{4} \right)^3}{\psi^{1/2}} \,d\psi = \int_{1}^{n^{\beta}}\!\!\! \frac{\sin \left(\psi - \frac{3 \pi}{4} \right)}{\psi^{1/2}}-\frac{\cos \left(\psi - \frac{3 \pi}{4} \right)^2 \sin \left(\psi - \frac{3 \pi}{4} \right) }{\psi^{1/2}} \,d\psi = \!\!\bigg[ \frac{\cos \left(\psi - \frac{3 \pi}{4} \right)}{\psi^{1/2}}\bigg|_{1}^{n^\beta}\\&- \int_{1}^{\pi n^\beta} \frac{\cos \left(\psi - \frac{3 \pi}{4} \right)}{\psi^{3/2}} \,d\psi  \bigg] -\bigg[ \frac{\cos \left(\psi - \frac{3 \pi}{4} \right)^3}{\psi^{1/2}}\bigg|_{1}^{n^\beta}- \int_{1}^{\pi n^\beta} \frac{\cos \left(\psi - \frac{3 \pi}{4} \right)^3}{\psi^{3/2}} \,d\psi  \bigg] < \infty;
	\end{split}
	\end{equation*}
	Similarly, for the other terms we find $O(\frac{1}{n^{\beta}}n^{\beta/2})$, $O(\frac{1}{n^{2\beta}} n^{3/2\beta})$ and $O(\frac{1}{n^{3\beta}} n^{5/2\beta})$.
	Putting together the results for $I_1$ and $I_2$ with (\ref{2var}) and (\ref{3var}) the thesis of the Theorem follows.
\end{proof}

\section{Proof of Lemma \protect\ref{Lemma1}}
Before proving Lemma \ref{Lemma1}, using the same notation as in \cite{FXA},
we recall the Hilb's asymptotic formula (see \cite{szego}, Theorem 8.21.12):
\begin{equation}  \label{hilbs}
P_n^{(1,0)}(\cos \theta ) = \bigg(\sin \frac{\theta}{2}\bigg)^{-1} \bigg\{ %
\bigg(\frac{\theta}{\sin \theta}\bigg)^{1/2} J_1((n+1)\theta)
+R_{1,n}(\theta) \bigg\}
\end{equation}
where $P_n^{(1,0)}$ is a Jacobi Polynomial, which in general is defined as 
\begin{equation*}
P_n^{(\alpha, \beta)}(x)= \sum_{s=0}^n \binom{n+\alpha}{s} \binom{n+\beta}{%
	n-s} \bigg(\frac{x-1}{2}\bigg)^{n-s}\bigg(\frac{x+1}{2}\bigg)^s,
\end{equation*}
\begin{equation}\label{error}
R_{1,n}(\theta) =%
\begin{cases}
\theta^3 O(n), & 0\leq \theta \leq c/n \\ 
\theta^{1/2} O(n^{-3/2}), & c/n \leq \theta \leq \pi -\epsilon%
\end{cases}%
\end{equation}
and $J_1$ is the Bessel function of order 1.
\begin{proof}[Proof of Lemma \ref{Lemma1}]
	Looking at the covariance function in (\ref{cov}) we can write $\bar{\Gamma}_{\alpha_{n,\beta}}(x,y)$ as \begin{equation}\label{var1}
	\begin{split}
	\bar{\Gamma}_{\alpha_{n,\beta}}(x,y)&= 
	C_{n,\beta} \bigg(\sum_{\ell=0}^n \frac{2\ell+1}{4\pi} P_{\ell}(\langle x,y \rangle)-\sum_{\ell=0}^{n\alpha_{n,\beta}-1} \frac{2\ell+1}{4\pi} P_{\ell}(\langle x,y \rangle )\bigg).
	\end{split}
	\end{equation}
	Thanks to the following formula (\cite{FXA}, page 6), derived by the Christoffel-Darboux formula (\cite{AH12}), 
	\begin{equation}\label{CDformula}
	\sum_{\ell=0}^{n} \sum_{m=-\ell}^{\ell} Y_{\ell, m} (x) Y_{\ell, m} (y)= \frac{n+1}{4\pi} P_n^{(0,1)}(\cos \theta(x,y)),
	\end{equation} 
	and to the addition formula (\cite{M e Peccati} page 66):
	\begin{equation}\label{addition formula}
	\sum_{m=-\ell}^{\ell} Y_{\ell, m} (x) Y_{\ell, m} (y)= \frac{2\ell+1}{4\pi} P_\ell(\cos \theta(x,y)),
	\end{equation}
	we obtain that 
	$$\bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)=C_{n,\beta} \bigg[ \frac{n+1}{4\pi} P_n^{(1,0)}(\cos \theta(x,y))-\frac{n\alpha_{n,\beta}}{4\pi} P_{n\alpha_{n,\beta}-1}^{(1,0)} (\cos \theta(x,y)) \bigg].$$
	Applying the Hilb's asymptotics (\ref{hilbs}) we get 
	\begin{equation}
	\begin{split}
	\bar{\Gamma}_{\alpha_{n,\beta}}(\cos \theta)=\frac{C_{n,\beta}}{4\pi}  \left(\sin \frac{\theta}{2}\right)^{-1}&\bigg[ (n+1) \left(\frac{\theta}{\sin \theta}\right)^{1/2} J_1((n+1)\theta)+(n+1)R_{1,n}(\theta)+\\&-n{\alpha_{n,\beta}} \left(\frac{\theta}{\sin \theta}\right)^{1/2} J_1(n \alpha_{n,\beta} \theta)- n\alpha_{n,\beta} R_{1,n\alpha_{n,\beta}}(\theta)  \bigg].
	\end{split}
	\end{equation}
	In view of (\ref{error}) the error term $(n+1)R_{1,n}(\theta)- n \alpha_{n,\beta} R_{1,n\alpha_{n,\beta}}(\theta) $,
	for $\dfrac{c}{n\alpha_{n,\beta}} \leq \theta \leq \pi-\epsilon$, it is equal to
	\begin{equation*}
	\theta^{\frac{1}{2}} O\!\left(n^{-\frac{3}{2}}\right)(n+1)-n \alpha_{n,\beta} \theta^{\frac{1}{2}} O\!\left(n^{-\frac{3}{2}} \alpha_{n,\beta}^{-3/2}\right)=\theta^{\frac{1}{2}} O\!\left(n^{-\frac{1}{2}}-n^{-\frac{1}{2}} \alpha_{n,\beta}^{-\frac{1}{2}}\right)= O\!\left(\frac{\theta^{\frac{1}{2}}}{\sqrt{n}}\right)=O\left(\frac{1}{\sqrt{n}}\right).
	\end{equation*}
	Now, changing variable $\theta= \frac{\psi}{\alpha_{n,\beta} n}$ and exploiting the expansion of the Bessel functions (\cite{szego}, page 15-16),
	\begin{equation}\label{J1}
	J_1 (x) = \bigg( \frac{2}{\pi x} \bigg)^{1/2} \cos \left(x - \frac{3\pi }{4}\right)- \dfrac{3}{4\sqrt{2\pi}x^{3/2}} \sin\left(x - \frac{3\pi }{4}\right)+O(x^{-5/2}), \mbox{ as } x \rightarrow \infty,
	\end{equation}
	we find
	\begin{equation*}
	\begin{split}
	&\bar{\Gamma}_{\alpha_{n}}\left(\cos \frac{\psi}{\alpha_{n,\beta} n}\right)=\frac{C_{n,\beta}}{4\pi}  \bigg(\sin \frac{\psi}{2\alpha_{n,\beta} n}\bigg)^{-1} \bigg\{ \left(\frac{\psi/(n\alpha_{n,\beta})}{\sin \psi/(\alpha_{n,\beta} n)}\right)^{1/2} \bigg[(n+1) \bigg( \bigg( \frac{2\alpha_{n,\beta} n}{\pi (n+1)\psi} \bigg)^{1/2} \\& \times\cos\! \left(\!\frac{\psi(n+1)}{\alpha_{n,\beta} n}-\!\!\frac{3\pi}{4}\right)-\frac{3}{4\sqrt{2\pi}} \left(\frac{1}{\psi}\right)^{3/2}\!\! \left(\frac{\alpha_{n,\beta} n}{n+1}\right)^{3/2}\! \!\!\sin \left(\frac{n+1}{\alpha_{n,\beta} n } \psi -\frac{3\pi}{4} \right)\!\!+\! O\left(\frac{\alpha_{n,\beta} n}{(n+1)\psi}\right)^{\frac{5}{2}} \!\!\bigg)\\&
	-n\alpha_{n,\beta} \bigg(\sqrt{\frac{2}{\pi \psi}} \cos\left(\psi-\frac{3\pi}{4}\right)- \frac{3}{4\sqrt{2\pi}} \left(\frac{1}{\psi}\right)^{3/2} \sin \left( \psi- \frac{3\pi}{4} \right) +O(1/\psi^{5/2}) \bigg) \bigg]+ O\left(\frac{1}{\sqrt{n}}\right)\bigg\}
	\end{split}
	\end{equation*}
	which is equal to
	\begin{equation*}
	\begin{split}
	&\bar{\Gamma}_{\alpha_{n}}\left(\cos \frac{\psi}{\alpha_{n,\beta} n}\right)
	=\frac{C_{n,\beta}}{4\pi}  \bigg(\sin \frac{\psi}{2\alpha_{n,\beta} n}\bigg)^{-1} \bigg\{ \left(\frac{\psi/(n\alpha_{n,\beta})}{\sin \psi/(\alpha_{n,\beta} n)}\right)^{1/2} \alpha_{n,\beta} n 
	\bigg[ 
	\sqrt{\frac{2}{\pi \psi}} \bigg( \bigg(\frac{n+1}{\alpha_{n,\beta} n} \bigg)^{1/2} \\& \times\cos \left((n+1)\frac{\psi}{\alpha_{n,\beta} n}-\frac{3\pi}{4}\right) - \cos \left( \psi- \frac{3\pi}{4} \right) \bigg)\\&
	+\frac{3}{4\sqrt{2\pi}} \left(\frac{1}{\psi}\right)^{3/2} \left( \sin\left(\psi-\frac{3\pi}{4}\right)-\left( \frac{\alpha_{n,\beta} n}{n+1}\right)^{1/2} \sin \left(\frac{n+1}{\alpha_{n,\beta} n } \psi -\frac{3\pi}{4} \right)\right)+ O\left(\frac{1}{\psi^{5/2}n^\beta} \right) \bigg]\bigg\}.
	\end{split}
	\end{equation*}
	Using the Taylor expansion 
	\begin{equation}\label{taylor}
	(1+x)^{\gamma}= 1+\gamma x+\frac{\gamma (\gamma-1)}{2}x^2+o(x^2) \mbox{ }\mbox{ }\mbox{ }\mbox{ }\mbox{ } (x \rightarrow 0)
	\end{equation} 
	applied to
	\begin{align*}
	\left(\frac{n+1}{\alpha_{n,\beta} n}\right)^{1/2}\!\!&=&\!\!\left(1+\frac{1}{n}\right)^{1/2}\left(1-\frac{1}{n^{\beta}}\right)^{-1/4}= 1+\frac{1}{4n^{\beta}}+\frac{5}{32n^{2\beta}}+\frac{1}{2n}+o\left(\frac{1}{n^{2\beta}}\right)+o\left(\frac{1}{n}\right),\\
	\left(\frac{\alpha_{n,\beta} n}{n+1}\right)^{1/2}\!\!&=&\!\!\left(1+\frac{1}{n}\right)^{-1/2}\left(1-\frac{1}{n^{\beta}}\right)^{1/4}= 1-\frac{1}{4n^{\beta}}-\frac{3}{32n^{2\beta}}-\frac{1}{2n}+o\left(\frac{1}{n^{2\beta}}\right)+o\left(\frac{1}{n}\right)
	\end{align*}
	we obtain
	\begin{equation}\label{uu}
	\begin{split}
	\bar{\Gamma}_{\alpha_{n}}&\left(\cos \frac{\psi}{\alpha_{n,\beta} n}\right)=\frac{C_{n,\beta}}{4\pi}  \bigg(\sin \frac{\psi}{2\alpha_{n,\beta} n}\bigg)^{-1} \bigg\{ \left(\frac{\psi/(n\alpha_{n,\beta})}{\sin \psi/(\alpha_{n,\beta} n)}\right)^{1/2} \alpha_{n,\beta} n 
	\bigg[ 
	\sqrt{\frac{2}{\pi \psi}} \\& \times\left( \cos \left(\frac{\psi(n+1)}{\alpha_{n,\beta} n}-\frac{3\pi}{4}\right) -  \cos \left( \psi- \frac{3\pi}{4} \right) \right) +\sqrt{\frac{2}{\pi \psi}} \frac{1}{4n^{\beta}} \cos \left((n+1)\frac{\psi}{\alpha_{n,\beta} n}-\frac{3\pi}{4}\right) \\&
	+\frac{3}{4\sqrt{2\pi}} \left(\frac{1}{\psi}\right)^{3/2} \left( \sin\left(\psi-\frac{3\pi}{4}\right)- \sin \left(\frac{n+1}{\alpha_{n,\beta} n } \psi -\frac{3\pi}{4} \right) \right)\\&+ \frac{3}{4\sqrt{2\pi}} \left(\frac{1}{\psi}\right)^{3/2} \frac{1}{4n^{\beta}}\sin\left(\psi\frac{n+1}{n\alpha_{n,\beta}}-\frac{3\pi}{4}\right)+O\left(\frac{1}{\sqrt{\psi}n^{2\beta}}+ \frac{1}{\psi^{3/2}n^{2\beta}}+\frac{1}{\psi^{5/2}n^\beta} \right) \bigg]\bigg\}.
	\end{split}
	\end{equation}
	Now for $n^{\beta} <\psi<\alpha_{n,\beta}n\pi$, using that
	\begin{eqnarray*}
		\cos \!\left(\frac{\psi(n+1)}{\alpha_{n,\beta} n}-\frac{3\pi}{4} \right) -  \cos \left( \psi- \frac{3\pi}{4} \right)\!\!&=&\!\!-2 \sin \left(	\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi\right) \sin \left(\frac{n+1}{n\alpha_{n,\beta}}\psi-\psi\right), \\
		\sin \!\left( \psi- \frac{3\pi}{4} \right)-\sin \left(\frac{\psi(n+1)}{\alpha_{n,\beta} n}-\frac{3\pi}{4} \right)\!\!&=&\!\!2 \cos \left(	\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi\right) \sin\left(\frac{\psi}{2}-\frac{n+1}{2n\alpha_{n,\beta}}\psi\right)
	\end{eqnarray*} we find
	\begin{equation}\label{gamma2}
	\begin{split}
	\bar{\Gamma}_{2}&=
	\frac{C_{n,\beta}}{4\pi}  \bigg(\sin \frac{\psi}{2\alpha_{n,\beta} n}\bigg)^{-1} \left(\frac{\psi/(n\alpha_{n,\beta})}{\sin \psi/(\alpha_{n,\beta} n)}\right)^{1/2} \frac{\alpha_{n,\beta} n}{\sqrt{\psi}} \\&\bigg[\sqrt{\frac{2}{\pi}} (-2) \sin \left(	\frac{\psi}{2}-\frac{3\pi}{4}+\frac{(n+1)}{2\alpha_{n,\beta} n} \psi\right) \sin\left(\frac{n+1}{n\alpha_{n,\beta}}\psi-\psi\right)+O\left( \frac{1}{ \psi} \right)\bigg].
	\end{split}%
	\end{equation}
	Whereas for $\psi \in [1,n^{\beta}] $ applying Taylor expansion to $\frac{n+1}{\alpha_{n,\beta}n}$, exploiting the formulas of sum and difference of angles and calling  $\zeta= \frac{\psi}{2n^{\beta}}+\frac{3\psi}{8n^{2\beta}} +\frac{\psi}{n}+o(\frac{1}{n})+o\left(\frac{1}{n^{2\beta}}\right)$ we have that 
	\begin{eqnarray*}
		\cos \left((n+1)\frac{\psi}{\alpha_{n,\beta} n}-\frac{3\pi}{4}\right)\!\!\!&=&\!\!\! \cos \left( \psi-\frac{3\pi}{4}+\zeta\right)=\cos \left(\psi-\frac{3\pi}{4}\right) \cos \zeta- \sin \left(\psi-\frac{3\pi}{4}\right) \sin \zeta,
		\\
		\sin \left((n+1)\frac{\psi}{\alpha_{n,\beta} n}-\frac{3\pi}{4}\right)\!\!\!&=&\!\!\! \sin \left( \psi-\frac{3\pi}{4}+\zeta\right)=\sin \left(\psi-\frac{3\pi}{4}\right) \cos \zeta+ \cos \left(\psi-\frac{3\pi}{4}\right) \sin \zeta.
	\end{eqnarray*} 
	Plugging these formulas in (\ref{uu}), 
	%
	using the fact that $\cos \zeta=1-\frac{\zeta^2}{2}+O(\zeta^4)$ and  $\sin \zeta=\zeta+ O(\zeta^3)$, as $\zeta \to 0$,
	we get 
	\begin{equation}\label{gamma1}
	\begin{split}
	\bar{\Gamma}_{1}&=
	\frac{C_{n,\beta}}{4\pi}  \bigg(\sin \frac{\psi}{2\alpha_{n,\beta} n}\bigg)^{-1} \bigg\{ \left(\frac{\psi/(n\alpha_{n,\beta})}{\sin \psi/(\alpha_{n,\beta} n)}\right)^{1/2} \alpha_{n,\beta} n 
	\bigg[ \sqrt{\frac{2}{\pi\psi}} \cos \left(\psi-\frac{3\pi}{4}\right)\\& (-1)\frac{\psi^2}{8n^{2\beta}}+O\left(\frac{1}{\sqrt{\psi}} \frac{\psi^2}{n^{3\beta}}\right)- \sqrt{\frac{2}{\pi\psi}} \zeta	\sin  \left(\psi-\frac{3\pi}{4}\right) +\sqrt{\frac{2}{\pi\psi}} \frac{1}{4n^{\beta}} \cos \left(\psi-\frac{3\pi}{4}\right)  \\&-\sqrt{\frac{2}{\pi}}\frac{1}{4n^\beta} \zeta \sin\left(\psi-\frac{3\pi}{4}\right)
	+\frac{3}{4 \sqrt{2\pi}} \frac{1}{\psi^{3/2}} 
	\bigg(\sin\left(\psi-\frac{3\pi}{4}\right)\frac{\psi^2}{8n^{2\beta}}+O\left(\frac{\psi^2}{n^{3\beta}}\right)  \\&-\zeta \cos \left(\psi-\frac{3\pi}{4}\right) +\frac{1}{4n^\beta}\sin \left(\psi-\frac{3\pi}{4}\right) \bigg)+O\left(\frac{\sqrt{\psi}}{n^{2\beta}}+ \frac{1}{\psi^{3/2} n^\beta}\right)+O\left(\frac{\psi^{3/2}}{n^{3\beta}}\right) \bigg]\bigg\}
	\\&=
	\frac{C_{n,\beta}}{4\pi}  \bigg(\sin \frac{\psi}{2\alpha_{n,\beta} n}\bigg)^{-1} \left(\frac{\psi/(n\alpha_{n,\beta})}{\sin \psi/(\alpha_{n,\beta} n)}\right)^{1/2} \alpha_{n,\beta} n \frac{\sqrt{\psi}}{n^\beta}\\& \times \sqrt{\frac{2}{\pi}} \frac{1}{2} \bigg(-\cos \left(\psi-\frac{3\pi}{4}\right) \frac{\psi}{4n^{\beta}}- \sin  \left(\psi-\frac{3\pi}{4}\right) +O\left( \frac{1}{ \psi} \right)\bigg) .
	\end{split}%
	\end{equation}
	Taylor expansion of $1/\sin x $ and $ \sin x$, as $x \to 0,$ implies that
	$$\frac{1}{\sin \psi/(n\alpha_{n,\beta})} \left( \frac{\psi/(\alpha_{n,\beta}n)}{\sin \psi/(n\alpha_{n,\beta})}\right)^{1/2}=\frac{\alpha_{n,\beta} n}{\psi}+  \frac{\psi}{\alpha_{n,\beta}6}+O\left(\frac{\psi^3}{(\alpha_{n,\beta} n )^3}\right)$$
	which replaced in (\ref{gamma1}) and (\ref{gamma2}) gives 
	the thesis of the Lemma.
\end{proof}

\subsection*{Acknowledgements}
This topic was suggested by Domenico Marinucci and Igor Wigman. The author
would like to thank them also for some useful discussions. Thanks also to an anonimous referee for some comments. The author has been supported by the German Research Foundation (DFG) via RTG 2131 and by GNAMPA-INdAM (project: \emph{Stime asintotiche: principi di invarianza e grandi deviazioni}).







\end{document}